OpenStudy (thatgirl21):

Will fan and give medal. Confused about how to solve this - If an average-sized man with a parachute jumps from an airplane, he will fall 12.5(0.2t − 1) + 19t feet in t seconds. How long will it take him to fall 142 feet? (Round your answer to two decimal places.) ____sec

OpenStudy (irishboy123):

calculus?

OpenStudy (thatgirl21):

Pre-calc for business

OpenStudy (thatgirl21):

so would i plug in 142 for v(t)?

OpenStudy (irishboy123):

yes!

OpenStudy (thatgirl21):

Oh also for the (0.2t) part it's supposed to be 0.2^t

OpenStudy (thatgirl21):

okay so it's 142 = 12.5(0.2^t-1)+19t. the first step is to distribute?

OpenStudy (irishboy123):

you can do all of that but it's ugly

OpenStudy (thatgirl21):

what i have so far is 154.5=2.5^t+19t. not sure if that's right

OpenStudy (irishboy123):

pre-calc for business? \(142 = 12.5(0.2^t-1)+19t \implies \color{red}{154.5 = 12.5\times 0.2^t+19t}\) there is no sensible way to solve that

OpenStudy (thatgirl21):

yes it's math 1643. I didn't think so

OpenStudy (irishboy123):

write some code, use some kinda numerical approach.

OpenStudy (irishboy123):

OU?

OpenStudy (thatgirl21):

yes

OpenStudy (irishboy123):

set t = 0 and you already get a shιt answer

OpenStudy (irishboy123):

case closed?